# Occlusion-Aware Ground Target Tracking by a Dubins Vehicle Using Visibility Volumes

### v1

@@ 9 Conclusion
This work describes an approach to track a ground POI moving along a known trajectory through an urban environment along a translating circular orbit with a time-varying radius that is inscribed inside the POI’s dynamic visibility volume. Visibility volumes are created by adaptively sampling points along the POI’s trajectory while satisfying a constraint that ensures the visibility do not change too rapidly. Conditions are derived that guarantee the resulting orbits are feasible for the Dubins vehicle. A series of linearly interpolated orbits are then created inside the discretized visibility to approximate a continuous visibility orbit. A Lyapunov-based control design is proposed to steer the UAV onto a vector field that guides the UAV onto the visibility orbit. The approach is demonstrated through several numerical examples and through a real-world multirotor UAV tracking experiment wherein a virtual POI is tracked in-between buildings and trees on a university campus. The flight test showed that the UAV mainted the POI within line-of-sight an average of 29% more often in comparison to a non-adaptive approach with a constant maximum-radius orbit for the chosen environment model.
Limitations of the proposed approach are that both the environment and POI motion are assumed to be known a priori. Future work may consider modifying the approach to allow for an estimated POI position and uncertain building geometry. Expanding the orbits from circles to arbitrary time-varying polar curves can also enable larger standoff distances that better approximate the visibility polygons.

### v2

@@ 8 Conclusion
This work describes an approach to track a ground POI moving along a known trajectory through an urban environment along a translating circular orbit with a time-varying radius that is inscribed inside the POI’s dynamic visibility volume (VV). VVs are created by adaptively discretizing points along the POI’s trajectory while satisfying a constraint that ensures the visibility do not change too rapidly. Conditions are derived that guarantee the resulting orbits are feasible for the Dubins vehicle. A series of linearly interpolated orbits are then created inside the discretized visibility to approximate a continuous visibility orbit. A Lyapunov-based control design is proposed to steer the UAV onto a vector field that guides the UAV onto the visibility orbit. The approach is demonstrated through several numerical examples and through a real-world multirotor UAV tracking experiment wherein a virtual POI is tracked in-between buildings and trees on a university campus. The flight test showed that the UAV maintained the POI within line-of-sight an average of 29% more often in comparison to a non-adaptive approach with a constant maximum-radius orbit for the chosen environment model. Limitations of the proposed approach are that both the environment and POI motion are assumed to be known a priori. Future work may consider modifying the approach to allow for an estimated POI position and uncertain building geometry. Expanding the orbits from circles to arbitrary time-varying polar curves can also enable larger standoff distances that better approximate the visibility polygons.